### v1

# Quantum Coherence of Topologically Frustrated Spin Chains

Abstract
The study of entanglement and magic properties in topologically frustrated systems suggests that, in the thermodynamic limit, these quantities decompose into two distinct contributions. One is determined by the specific nature of the model and its Hamiltonian, and another arises from topological frustration itself, resulting in being independent of the Hamiltonian’s parameters. In this work, we test the generality of this picture by investigating an additional quantum resource, namely quantum coherence, in two different models where topological frustration is induced through an appropriate choice of boundary conditions. Our findings reveal a perfect analogy between the behavior of quantum coherence and that of other quantum resources, particularly magic, providing further evidence in support of the universality of this picture and the topological nature of this source of frustration.
preprint: RBI-ThPhys-2025-19
I Introduction
Although many-body quantum mechanics is known to exhibit a significantly richer phenomenology than its classical counterpart, the application of concepts borrowed from classical statistical mechanics has led to notable successes. Thus, following Landau’s approach, phases have been classified based on quantities such as correlation functions and local order parameters [1]. However, the breadth of the phenomenology of many-body quantum systems is too broad to be fully described by such an approach. A prominent example of a phenomenon that eludes such a classification is the existence of topological phases. Such phases are characterized by discrete-valued nonlocal quantities that remain invariant under continuous deformations in the parameter space [2]. This phenomenon can be understood, for instance, in terms of the presence of one or more topological defects, which can be localized at the edges, such as the Majorana zero modes in the Kitaev wire [3], or can propagate freely in the bulk in the form of topological solitons, as in the fractional quantum Hall effect [4, 5].
Since their discovery, several models have shown phenomenologies, often very different from each other, which can be encapsulated into the class of topological effects. Among them, the topologically frustrated (TF) models have attracted considerable interest in recent years [6, 7, 8]. This phenomenon was first identified in one-dimensional antiferromagnetic spin chains (AFMs) with nearest-neighbor coupling and frustrated boundary conditions (FBCs) [9, 10]. FBCs occur when periodic boundary conditions (PBCs) are imposed on a system with an odd number of spins. This configuration breaks the periodicity of the AFM interaction, introducing a delocalized excitation into the system. As a result of the presence of this excitation, the static [11, 12] and dynamical [13] properties of TF systems acquire distinctive and universal behaviors. These results aroused considerable interest, which further increased when it was shown that the phenomenon of topological frustration extends beyond the class of models in which it was initially unveiled. A notable example is the so-called ANNNI model [14], characterized by competing interactions between couples of spins at nearest and next-nearest neighbor sites along the lattice. In this model, topological frustration occurs when a system with an even number of spins, not divisible by four, is subjected to periodic boundary conditions, resulting in the emergence of two delocalized excitations.
Among the various effects of imposing topological frustration on a one-dimensional system, one of the most surprising involves two fundamental quantum resources [15] such as the entanglement [16] and the non-stabilizerness (also known as magic) [17]. Specifically, in the frustrated phase [10, 18, 6, 19], as the chain length diverges, the values of these resources can be decomposed into two distinct contributions. The first is a somewhat local term, corresponding to the value the resource would assume when topological frustration is removed, i.e. when a different set of boundary conditions is applied to the system. In contrast, the second term has a topological nature and, therefore, remains independent of the specific Hamiltonian parameters. Thus, its value, which remains constant throughout the phase, can be evaluated (analytically) near the classical point, namely the point at which the Hamiltonian reduces to a sum of mutually commuting terms and the ground state can be calculated perturbatively. This contribution to the quantum resources can be attributed to the presence of delocalized excitations induced by TF, which are known to be responsible for the topological nature of the TF phase, as proved through the Disconnected Entanglement entropy (DEE) [20]. Indeed, a finite DDE, such as that found in [20], is a unique signature of the long-range correlation typical of topological phases, although a characteristic topological invariant has not been identified yet for TF.
It is therefore natural to ask how general this picture is and whether or not it extends to other quantum resources. Unfortunately, various quantum resources can have different structures with few points of contact, and thus it does not seem feasible to provide a general proof of this property or to perform a broad analysis. For this reason, accepting our limits, in this paper we test the hypothesis of the generality of this picture by focusing on a further resource: Quantum Coherence (QC) [21, 22] The choice of the QC is not accidental but takes into account its relevance in several fields of research. Indeed, QC plays a fundamental role across multiple areas of quantum physics. In quantum information, QC is a key resource for quantum computation and communication, enabling quantum speedups and secure protocols [21, 23]. In quantum thermodynamics, QC influences work extraction, quantum fluctuation theorems, and the fundamental limits of energy transfer in nanoscale systems [24, 25, 26, 27]. In quantum many-body systems, its analysis can underlie the presence of quantum phase transitions, and the emergence of macroscopic quantum phenomena [28, 22, 29, 30]. Moreover, QC is essential for understanding the interplay between entanglement and non-classical correlations in open quantum systems [31].
To achieve the objective of this paper, it is necessary to disentangle the two contributions to QC. The topological contributions to QC are derived purely analytically by considering the system in proximity of the classical point. All other data are obtained through numerical computations. Given the exponential complexity of the evaluation, we employ the Density Matrix Renormalization Group (DMRG) algorithm to efficiently represent the ground state of the system and the tensor-cross-interpolation (TCI) algorithm to construct an efficient representation of QC as a univariate function [32]. Our results indicate that as the system size increases, the QC in the frustrated phase becomes increasingly well approximated by the sum of the topological and the local terms. This observation, together with previous findings for entanglement entropy and stabilizer Rényi entropy, suggests a universal behavior for at least all entropy-based quantum resource quantifiers.
The paper is structured as follows. In Sec.II, we provide a brief overview of the numerical approach proposed in [32] for computing the QC as measured by the Relative Entropy of Coherence (REC). In Sec. III, we describe the models under consideration and summarize their known phenomenology. Section IV presents our findings, demonstrating that, similar to entanglement and magic, REC decomposes into two distinct contributions: a local term, corresponding to the REC value in the absence of TF, and a second contribution of topological origin. Finally, in Sec. V we summarize the results and discuss future developments, and in Appendix A we provide the analytical calculation of REC for the ground state of the ANNNI chain.
II Measuring Quantum Coherence
Over the years, several different measures of QC have been suggested [22], each of them with its positive and negative sides. Since we are interested solely in pure states, in this work, we resort to the Relative Entropy of Coherence (REC) [21]. For a state with a density matrix $\rho$ the REC is defined as $C(\rho)=S(\rho_{\textrm{diag}})-S(\rho)$, where $S(\rho)$ stands for the von Neumann entropy and $\rho_{\textrm{diag}}$ is the matrix obtained by $\rho$ by deciding on a reference basis frame and setting all off-diagonal elements in this basis to zero. For a pure state, i.e. for $\rho=\ket{\psi}\!\bra{\psi}$ |, $S(\rho)$ vanishes and the REC reduces to
$C(\rho)=-\sum_{i=1}^{2^{L}}c_{i}^{2}\log_{2}c_{i}^{2}.$ (1)
where $L$ is the number of qubits in the system, and $c_{i}$ are the coefficients of the state |$\ket{\psi}$, in the base {|$\{\ket{i}\}_{i=1}^{2^{L}}$, i.e. |$\ket{\psi}=\sum_{i=1}^{2^{L}}c_{i}\ket{i}$. It is worth emphasizing that the expression for REC in (1) clearly shows that, like all other QC measures, this quantity depends on the chosen basis for calculation. Throughout this work, we will consistently evaluate REC on the computational basis.
Despite its apparent simplicity, the computation of Eq. (1) is challenging, and, for large systems, it becomes unfeasible without an efficient representation of |$\ket{g}$. But luckily this kind of representation exists and it is provided by the Matrix Product States (MPS) [33, 34, 35, 36, 37]. Within such a representation, to evaluate the REC we observe that it is the sum of the components of the vector $f(\boldsymbol{c}^{\intercal})=(f(c_{1}),\ldots,f(c_{n}))$, obtained by applying the function $f(x):=-x^{2}\log_{2}x^{2}$ element-wise to the system’s ground state. This sum can be evaluated by sampling the set $\{f(c_{i})\}_{i=1}^{2^{L}}$ exploiting the family of tensor cross-interpolation (TCI) algorithms [38, 39] that allows us to escape the exponential scaling of the number of coefficients with the size of the system $L$ in (1[40]. It is worth noticing that, TCI algorithms are stable; within a given tolerance, the system consistently converges after a sufficient number of function evaluations for a specified bond dimension [38]. To summarize, to evaluate the REC we follow a three-step algorithm [32]. At first, we use a DMRG code to compute the ground state of the system. Hence, we exploit a TCI algorithm to sample the function $f$ of the ground state coefficients. Finally, to evaluate the REC, we contract the MPS resulting from the TCI decomposition.
III Models
Using REC, we tested our hypothesis on quantum resources on two different one-dimensional models that can, under suitable conditions, show the presence of TF. Both these models can be obtained for a particular choice of parameters, from the following Hamiltonian
$H=J_{1}\sum_{i=1}^{L}\sigma_{i}^{z}\sigma_{i+1}^{z}+J_{2}\sum_{i=1}^{L}\sigma_{i}^{z}\sigma_{i+2}^{z}+h\sum_{i=1}^{L}\sigma_{i}^{x}.$ (2)
In Eq. (2) $\sigma_{i}^{\alpha}$ with $\alpha=x,\,y,\,z$ stand for the Pauli operators acting on the $i$-th spin of the system, $J_{1}$ and $J_{2}$ are, respectively, the next-neighbor and the next-to-nearest neighbor interactions, and $h$ is a local transverse field. Unless stated otherwise, we assume periodic boundary conditions, i.e. $\sigma_{i}^{\alpha}\equiv\sigma_{i+L}^{\alpha}$.
Setting $J_{2}=0$ we recover the well-known Ising model with a transverse field [41, 42] that shows a quantum phase transition at $|h|=h_{c}\equiv 1$. Choosing $J_{1}>0$ and imposing FBCs, for $0<|h|<h_{c}$, the system is in a topologically frustrated phase with a gapless energy spectrum in the thermodynamic limit. For $h=0$ (the classical point mentioned above), there are $2L$ exactly degenerate lowest energy states, namely the mutually orthogonal kink states |k+$\ket{k^{+}}$ and |$\ket{k^{-}}$, which are individually separable and diagonal in the computational basis. Indeed, because of PBCs and $L$ being odd, it is impossible to have perfect Neel (staggered) order these states present a magnetic defect, i.e. two parallel aligned neighboring spins, which can be placed on any bond of the ring. Hence |k+$\ket{k^{+}}$ and |$\ket{k^{-}}$ are the two states with the defect between the $k$-th and the $k+1$-th sites and differ from each other by the orientation of the defect. A finite magnetic field hybridizes these states and lifts the degeneracy, resulting in the aforementioned energy band. In particular, in the limit $h\rightarrow 0^{+}$ the GS is given by |$\ket{g}=\frac{1}{\sqrt{2L}}\sum_{k=1}^{L}(\ket{k^{+}}+\ket{k^{-}})$ and the whole band can be characterized as a single delocalized kink excitation with some momentum.
On the contrary, considering $J_{2}>0$, we recover the so-called ANNNI model in which, independently of boundary conditions and the sign of $J_{1}$, there is always geometrical frustration induced by the competition between the two interactions. The phase diagram of this model is rich, having four distinct phases depending on the dominant interaction and the value of the magnetic field [14]. Among these, one of the most peculiar is the so-called antiphase which can occur when the next-to-near neighbor interaction $J_{2}$ dominates over $J_{1}$, i.e. if and only if $\kappa=|J_{1}|/J_{2}>1/2$. In such a phase, TF can be induced considering a ring made of an even number of spins that is not an integer multiple of four ($L=4n+2,n\in\mathbb{N}$). On the classical line $h=0$, the presence of TF increases the degeneration of the ground state manifold, which goes from a dimension equal to 4 (in the absence of TF) to $L^{2}/2$ if $J_{1}\neq 0$ or to $L^{2}$ in the limiting case $J_{1}=0$, where the chain decomposes into two independent subchains for the even and odd sites. Turning on a small transverse field, the degeneracy is lifted and the system admits a unique ground state given in Eq. (4), that can be described in terms of a distance-modulated superposition of kink states [19] (see Appendix A for details). Regardless of the presence or absence of the TF, by fixing the value of the ratio $\kappa$ and further increasing the value of $h$, the system undergoes two successive phase transitions. In the first, the system exits the antiphase and enters the so-called floating phase, whose nature has not yet been fully understood. It is characterized by a violation of the area law in terms of entanglement and is conjectured to be in a Luttinger liquid phase with algebraic correlations. Subsequently, the second transition marks the passage from the floating phase to a paramagnetic phase, where the dynamic is dominated by the transverse field.
Figure 1: Upper Panel: Quantum Coherence of the ground state of the TF Ising chain (red full markers) and of the corresponding unfrustarted FM model (blue empty markers) as a function of the magnetic field and for different values of the system size. Middle Panel: The same quantities are plotted as a function of the system length for different values of the magnetic field. Bottom Panel: The ratio $R$ in eq. (3) for the Ising model as a function of the magnetic field and for different system sizes.
IV Results
Having introduced the models of interest, let us discuss our results for the calculation of REC on the computational basis. It is crucial to emphasize that the choice of basis for calculating QC is fundamental, as QC, like magic but unlike entanglement, is a basis-dependent resource. Among various options, we choose to work on the computational basis because, near the classical point and in the absence of topological frustration, quantum coherence reaches its minimum.
In Fig.1, we present our findings for the Ising model $(J_{2}=0)$. In the upper panel, we display the raw data obtained for the REC as a function of the applied magnetic field, both with $(J_{1}=1)$ and without $(J_{1}=-1)$ TF. Although our data are derived from systems of finite sizes, two distinct behaviors corresponding to the two phases of the model are observable. For large magnetic field $h$, the REC of the TF and non-TF Ising model converges to the same value. On the other hand, when the interaction term dominates over the external field, the behaviors differ significantly, and the difference increases with the system size. In particular, the limit $h\rightarrow 0^{+}$ is of special interest in this region. For the unfrustrated system, as $h$ approaches zero, $C(\rho)$ becomes independent of $L$ and tends to 1, consistent with the fact that the ground state can be well approximated by the GHZ state |$\ket{\phi}=\frac{1}{\sqrt{2}}(\ket{\uparrow}^{\otimes L}+\ket{\downarrow}^{\otimes L})$, where $\ket{\uparrow}$ and $\ket{\downarrow}$ are the eigenstates of $\sigma^{z}$ with eigenvalues $+1$ and $-1$, respectively. In contrast, with TF, for $h\to 0^{+}$ the REC tends to $1+\log_{2}(L)$, consistently with what can be inferred from the expression for the ground state of the TF Ising model presented in the previous section.
Increasing $h$ but remaining in the same phase, the independence on the size disappears and the unfrustrated system shows a volume law, which recalls the behavior of magic, as measured by the stabilizerness Rényi entropy [6], as can be seen in the middle panel of Fig. 1. The analogy between the stabilizer Rényi entropy and the REC holds in TF systems as well, where both exhibit a volume-law behavior in the large-size limit, although with a significant correction for smaller systems.
To test our hypothesis that, in the thermodynamic limit, any quantum resource evaluated for the ground state of a TF model becomes the sum of a local term and a topological one that does not depend on the Hamiltonian parameters, we adopt the same approach as used in [18] for the Entanglement Entropy and in [6, 8] for the Stabilizer R’enyi Entropy. This approach involves the study of the ratio
$R=\frac{Q(\rho_{\textrm{fr}})-Q(\rho_{\textrm{unfr}})}{Q(\rho_{\textrm{fr}}^{h\rightarrow 0^{+}})-Q(\rho_{\textrm{unfr}}^{h\rightarrow 0^{+}})},$ (3)
where $Q$ represents an entropic measure of a generic quantum resource, while $\rho_{\textrm{fr}}$ and $\rho_{\textrm{unfr}}$ denote the density matrices of the ground states for the topologically frustrated and topologically unfrustrated models, respectively, given the same Hamiltonian parameters.
If our initial hypothesis holds, the ratio in Eq. (3) should converge to 1 in the frustrated phase and approach zero in the paramagnetic phase. To compute this ratio, we first need to determine the denominator. For the Ising model, as discussed in the previous section, near the classical point and in the presence of topological frustration, the ground state can be approximated as a homogeneous superposition of $2L$ elements from the computational basis, yielding a REC of $\log_{2}(2L)$. Conversely, in the same region of parameter space but without frustration, the ground state simplifies to a GHZ state with a REC of 1. Thus, the denominator of Eq. (3) reduces to $\log_{2}(L)$. The bottom panel of Fig.1 shows our numerical results for $R$ in the case of the Ising model as a function of the magnetic field. Consistent with our working hypothesis, we observe that, as $L$ diverges, the value approaches $1$ throughout the whole topological phase and zero in the paramagnetic phase. The speed at which these values are reached depends on the ratio between the correlation length and the system size. Since the correlation length diverges at the phase transition, finite-size effects vanish more rapidly the further one moves away from $h=1$. We can then write the REC in the thermodynamic limit as the sum of the topological term and the magnetic one.
Figure 2: Upper Panel: Quantum Coherence of the ground state of the TF ANNNI chain (red full markers) and of the corresponding unfrustarted OBCs model (blue empty markers) as a function of the magnetic field and for different values of the system size. Middle Panel: The same quantities are plotted as a function of the system length for different values of the magnetic field. Bottom Panel: The ratio $R$ in eq. (3) for the ANNNI model as a function of the magnetic field and for different system sizes.
Let us now consider the ANNNI model by introducing an antiferromagnetic next-to-nearest-neighbor interaction ($J_{2}>0$). Even if the two models seem not so different from each other, the ANNNI model is much more complex to analyze, and several technical challenges must be addressed. The first problem arises when trying to associate the specific non-TF model with its TF equivalent. For the Ising model, this link is natural. Indeed, assuming periodic boundary conditions (PBCs) and an odd number of spins, transitioning between TF and non-TF Ising systems can be achieved simply by inverting the sign of $J_{1}$. This approach does not apply to ANNNI models. Specifically, when PBCs are imposed and the number of spins is even but not an integer multiple of 4, the presence of the TF is unaffected by the sign of $J_{1}$. On the other hand, by changing the sign of $J_{2}$, we not only remove TF but also the extensive frustration characteristic of the model, thereby fundamentally altering the properties of the system. Therefore, according to [19], this comparison is done by opening up the ring and moving from PBCs to open boundary conditions (OBCs). This introduces finite-size effects in the unfrustrated system which are expected to become irrelevant in the thermodynamic limit.
The results for REC in the ANNNI model are shown in Fig. 2. In analogy with the Ising model, the presence of TF induces, in the limit $h\rightarrow 0^{+}$, a dependence of REC on $L$ that is directly related to the dimension of the ground state manifold at the classical point. In this case, it is possible to show (see Appendix A) that, in the limit $h\rightarrow 0^{+}$, the REC of the TF system becomes equal to $C(\rho_{\textrm{fr}}^{h\rightarrow 0^{+}})=\log_{2}[L(L+2)]-\log_{2}(e)$. On the contrary, without topological frustration, the system allows a ground state that is a homogeneous superposition of four elements of the computational basis, and hence $C(\rho_{\textrm{unfr}}^{h\rightarrow 0^{+}})=2$. As $h$ increases, the difference between the REC for the TF and non-TF models decreases, until, in the floating phase, the REC for the non-TF model becomes larger than that for the TF model. This behavior appears to be a characteristic feature of the floating phase and could serve as a starting point for future investigations of this elusive phase of the model.
The lower panel of Fig.2 shows the behavior of the ratio in (3) as a function of $h$ for a fixed value of $\kappa=|J_{1}|/J_{2}>1/2$ and allows us to test our working hypothesis. As we can see, despite the relatively small size of the system considered, constrained by the large bond size in the DMRG algorithm required to obtain an accurate estimate of the ground state, the thermodynamic behavior is evident. In the TF phase, the ratio Eq. (3) approaches 1 while, in the paramagnetic phase, it drops to 0 in a complete analogy with the Ising model.
V Conclusions and outlook
In summary, we examined quantum coherence in two one-dimensional models—the Ising and the ANNNI model—comparing its behavior in the presence and absence of topological frustration. In both cases, we found that in the paramagnetic phase, quantum coherence remains unaffected by this source of frustration in the thermodynamic limit. However, in the frustrated phase, topological frustration—though not the only source of frustration in the ANNNI model—introduces a correction that scales logarithmically with the system size.
Our findings confirm that, at least in one-dimensional models, topological frustration induces in quantum coherence the same structural behavior previously identified in entanglement and magic. This leads to a clear separation between two contributions: a local term dependent on the Hamiltonian parameters and a topological term that remains constant throughout the phase.
Leveraging its phase invariance, and drawing parallels with other quantum resources, we analytically derived the expression for the contribution associated with topological frustration using the explicit ground state near the classical point. This separation is further emphasized by the distinct scaling behaviors of the two terms: for quantum coherence, as in magic [6], the local contribution scales extensively, while the topological component follows a logarithmic dependence on system size, reflecting the dimension of the ground state manifold at the classical point. In contrast, for entanglement, the local term follows an area law, remaining independent of system size $L$, whereas the topological term depends on the logarithm of the ratio between the length of the subsystem $M$ and $L$.
At this point, a natural question arises: should the behavior observed for quantum coherence, entanglement, and magic be regarded as a universal feature of all quantum resources when evaluated on the ground states of topologically frustrated systems in the thermodynamic limit? While identifying multiple cases where this pattern holds does not constitute a formal proof, the fact that all analyzed quantum resources exhibit the same behavior strongly supports this conjecture.
[…]
The analysis carried out in this work was made possible by using the algorithm introduced in [32] to calculate the relative entropy of coherence. This method leverages the efficient MPS representation of the system’s ground state in combination with the TCI algorithm class to perform measurements. By adopting this approach, it is possible to overcome the challenge of the exponential growth of the system’s Hilbert space, hence exploring larger system sizes. Furthermore, this technique is sufficiently general to compute a wide range of observables and quantum resources, offering the potential to further investigate the properties of topologically frustrated systems in the thermodynamic limit, a direction we plan to pursue in future work.
Acknowledgements.
SBK and GT thank Nora Reinić, Daniel Jaschke, and Simone Montangero for useful discussions and valuable insights. SBK acknowledges support from the Croatian Science Foundation (HrZZ) Projects DOK-2020-01-9938.
Appendix A Quantum Coherence of ANNNI ground state near the classical point
In this Appendix, we compute analytically the quantum coherence of the ground state of the TF ANNNI chain made of a number of spins that is even but not an integer multiple of four near the classical point. To do this, we will exploit the ground state expression achieved in [19] using perturbation (and graph) theory valid for $h\ll 1$
|$\ket{g}=A\sum_{k=1}^{L}\sum_{p=k}^{k+L/2-1}\!\!\sin\left[\frac{(p-k+1)L\pi}{L+2}\right]\ket{\psi(k,p)}.$ (4)
[…]
Applying eq. (1) to the state in eq. (4), the quantum coherence becomes
$C(\rho_{\textrm{fr}}^{h\rightarrow 0^{+}})$ $=$ $-\frac{4}{L+2}\sum_{r=0}^{\frac{L}{2}-1}\sin^{2}\left[\alpha(r+1)\right]\crossproduct$ (5)
×$~{}~{}~{}\crossproduct\log_{2}\left[\frac{4}{L(L+2)}\sin^{2}\left[\alpha(r+1)\right]\right]\text{,}~{}~{}~{}~{}$
where $\alpha=L\pi/(L+2)$. To obtain a closed form, we start splitting Eq. (5) in two terms separating the $L$ contribution in the logarithm from the $\sin$ function. The first term is given by
$\frac{-4}{L+2}\log_{2}\!\left[\frac{4}{L(L\!+\!2)}\right]\!\!\sum_{r=0}^{\frac{L}{2}-1}\sin^{2}\left[\alpha(r\!+\!1)\right]=\log_{2}\!\left[\frac{L(L\!+\!2)}{4}\right],$
where we have explicitly taken into account that the number of sites is even. The second term does not admit a simple closed form. To handle it we consider its thermodynamic limit ($L\rightarrow\infty$)
$\frac{-4}{L+2}\int_{-\frac{1}{2}}^{\frac{L}{2}+\frac{1}{2}}\mathrm{d}x\sin^{2}\left[\alpha(x+1)\right]\log_{2}\left[\sin^{2}\left[\alpha(x+1)\right]\right]=$
$\frac{-4}{\ln(2)(L+2)}\int_{-\frac{1}{2}}^{\frac{L}{2}+\frac{1}{2}}\mathrm{d}x\sin^{2}\left[\alpha(x+1)\right]\ln\left[\sin^{2}\left[\alpha(x+1)\right]\right]=$
$=\frac{-4}{L+2}\left(\frac{L+2}{2}\right)\frac{1}{\ln(2)}\times$
×$\times\left(\frac{1}{\frac{\alpha(L+2)}{2}}\int_{\frac{\alpha}{2}}^{\frac{\alpha}{2}(L+3)}\mathrm{d}y\sin^{2}(y)\ln\left[\sin^{2}(y)\right]\right)$
$=\frac{1}{\ln(2)}(2\ln(2)-1)$
where we averaged the integral in parenthesis over a long interval, which we can safely replace with an average over the period $[0,2\pi]$. Combining the two contributions, we finally find in the $L\rightarrow\infty$ limit
$C(\rho)=\log_{2}[L(L+2)]-\log_{2}(e).$ (6)
In Fig. 3, we compare this formula with the numerical results obtained through the procedure outlined in Section II, finding good agreement even for small systems. To further support our claim, we also compare our results with the numerics derived from the exact expression for the ANNNI ground state, given by Eq.(4), near the classical point [19].
Figure 3: Comparison of the different results for the REC close to the classical point of the ANNNI model. The black line stands for the analytic expression in Eq. 6, a The blue circular points indicate the results obtained from the exact diagonalization of the ANNNI in the first-order perturbation theory ($h\rightarrow 0^{+}$). The red squares indicate the results of a numerical analysis that exploit the procedure outlined in Sec. II with Hamiltonian parameters equal to $h=10^{-}3$ $\kappa=1.0$ and tolerance set equal to $\epsilon=10^{-12}$